OpenStudy (anonymous):

If the equation y = x + 4 is changed to y = x - 2, how will the graph of the equation change?

OpenStudy (anonymous):

Since the slope is the same, but the y-intercept changed from 4 to -2, this is a transposition of -6. In other words, the graph moved down by 6 units.

OpenStudy (anonymous):

if your function is y=f(x) and you are changing it by y=f(x)+a then graph of y=f(x) will moved up by 'a' unit... if your eq is y=f(x+a) then your graph will move 'a' unit leftwardly... in this case you have added -6 in y=f(x)..hence it will move up by -6 unit means it will move down by 6 unit.